Van Doren Housing expects to have sales this year of $15 million under its current credit policy. The present terms are net 30; the days sales outstanding (DSO) is 60 days; and the bad debt loss percentage is 5 percent. Also, Van Doren’s cost of capital is 15 percent, and its variable costs total 60 percent of sales. Since Van Doren wants to improve its profitability, a proposal has been made to offer a 2 percent discount for payment within 10 days; that is, change the credit terms to 2/10, net 30. The consultants predict that sales would increase by $500,000, and that 50 percent of all customers would take the discount. The new DSO would be 30 days, and the bad debt loss percentage on all sales would fall to 4 percent.

What is: the cost of the discounts taken; the incremental bad debt losses if the change were made; and, the incremental cost of carrying if the change were made?
Answer #1

1)

Cost of discount = ((sales +increase in sales) * Discount rate * probability of customer

= ((15000000+50000)*2%*50%)

= 155000

2)

Incremental bad debts = (((sales +increase in sales) * percent of bad debts) – ( sales * percent of bad debts)

= (((15000000+500000)*4%)-(15000000*5%))

=130000.

3)

At sales of 15000000 = (DSO * (sales/days)* percent of sales * cost of capital)

= (60*(15000000/360)*60%*15%)

= 225000

At sales of 15500000 = (DSO * (sales/days)* percent of sales * cost of capital)

= (30*(15500000/360)*60%*15%)

= 116250

Incremental cost of carrying =116250-225000 =108750.